Lott Company uses a job order cost system and applies overhead to production on the basis of direct labor costs. On January 1, 2017, Job No. 50 was the only job in process. The costs incurred prior to January 1 on this job were as follows: direct materials $23,200, direct labor $13,920, and manufacturing overhead $18,560. As of January 1, Job No. 49 had been completed at a cost of $104,400 and was part of finished goods inventory. There was a $17,400 balance in the Raw Materials Inventory account. During the month of January, Lott Company began production on Jobs 51 and 52, and completed Jobs 50 and 51. Jobs 49 and 50 were also sold on account during the month for $141,520 and $183,280 respectively. The following additional events occurred during the month. 1. 2. 3. Purchased additional raw materials of $104,400 on account. Incurred factory labor costs of $81,200. Of this amount $18,560 related to employer payroll taxes. Incurred manufacturing overhead costs as follows: indirect materials $19,720; indirect labor $23,200; depreciation expense on equipment $13,920; and various other manufacturing overhead costs on account $18,560. Assigned direct materials and direct labor to jobs as follows. 4. Job No Direct Materials Direct Labor $5,800 29,000 23,200 50 $11,600 45,240 34,800 52 X Your answer is incorrect. Try again. Calculate the predetermined overhead rate for 2017, assuming Lott Company estimates total manufacturing overhead costs of $974,400, direct labor costs of $812,000, and direct labor hours of 23,200 for the year. (Round answer to the nearest whole percent, e.g. 25%.) Predetermined overhead rateOpen job cost sheets for Jobs 50, 51, and 52. Enter the January 1 balances on the job cost sheet for Job No. 50 Job No. 50 Date Beg Direct Materials Direct Labor Manufacturing Overhead 23200 13920 18560 Jan 11600 5800 6960 37800 19720 25520 Cost of completed job Direct materials 34800 Direct labor 19720 Manufacturing overhead 25520 Total cost 80040 Job No. 51 Date Direct Materials Direct Labor Manufacturing Overhead Jan 45240 29000 34800 45240 29000 37800 Cost of completed job Direct materials 4524019720 25520 Cost of completed job Direct materials 34800 Direct labor 19720 25520 Total cost 80040 Job No. 51 Date Direct Materials Direct Labor Manufacturing Overhead Jan. 45240 29000 34800 45240 29000 Cost of completed job Direct materials 45240 Direct labor 29000 34800 Total cost 109040 Job No. 52 Date Direct Materials Direct Labor Manufacturing Overhead Jan.Prepare the journal entries to record the purchase of raw materials, the factory labor costs incurred, and the manufacturing overhead costs incurred during the month of January. (Credit account titles are automatically indented when amount is entered. Do not indent manually.) No. Account Titles and Explanation (0 Debit CreditPrepare the journal entries to record the assignment of direct materials, direct labor, and manufacturing overhead costs to production. In assigning manufacturing overhead costs, use the overhead rate calculated in (a). (Credit account titles are automatically indented when amount is entered. Do not indent manually.) No. Account Titles and Explanation Debit CreditTotal the job cost sheets for any job(s) completed during the month. Prepare the journal entry to record the completion of any job(s) during the month. (Credit account titles are automatically indented when amount is entered. Do not indont manually.) Account Titles and Explanation Debit Credit LIST OF ACCOUNTS Your answer is incorrect. Try again. Prepare the journal entries to record the sale of any job(s) during the month. (Credit account titles are automatically indented when amount is entered. Do not indent manually.) No. Account Titles and Explanation Debit Credit (To record sale of jobs) To record cost of joba)What is the balance in the Finished Goods Inventory account at the end of the month? What does this balance consist of? Finished Goods Inventory SHOW LIST OF ACCOUN Your answer is ncorrect. Try again. What is the amount of over- or underapplied overhead? Manufacturing Overhead
Answer #1

Predetermined overhead rate = manufacturing overhead cost / direct labour cost

=$974400/812000=120%

In job 50,51

Where the incorrect amount of $37800 is marked

Correct amount is $34800

= 23200 +11600=$34800

In job 52

Direct material =$34800

Direct labour=$23200

Manufacturing overhead =23200*120%=$27840

Journal entry for purchase of raw materials

Raw material a/c Dr. 104400

To supplier a/c 104400

Jounal entry for direct labour exp

Direct labour a/c Dr. 81200

To Cash 62640

To payroll taxes a/c payable. 18560

Journal entry for manufacturing overhead incurred

Indirect materials a/c dr 19720

Indirect labour a/c dr. 23200

Depreciation a/c dr. 13920

Other overheads a/c Dr. 18560

To cash a/c 75400

Journal entries for assigning manufacturing costs

Job 50

Work in process inventory a/c Dr 80040

To manufacturing exp. 80040

Job 51

Work in process inventory a/c dr 109040

To manufacturing exp. 109040

Job 52

Work in process inventory a/c dr 85840

To manufacturing exp. 85840

For completion of job

Finished good a/c dr 104400

To work in process a/c 104400

For sale of job 49

Debtor a/c dr 141520

To sales a/c 141520

To record cost of job 49

Cost of production a/c dr 104400

To finished goods inventory a/c 104400

Finished goods inventory account consists of inventory from jobs= 80040+109040=189080

For amount of over /under applied manufacturing overhead

Actual manufacturing overhead= $75400

Applied overheads =120%* 58000=$69600

Underapplied overhead =$5800

Do give your feedback ! Happy learning :)